40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 1 20:20:15 2016, gautam, Update, IMC, IMC loss measurement plan IMG_3471.JPG
    Reply  Fri Dec 2 16:40:29 2016, gautam, Update, IMC, 24V fuse pulled out 
       Reply  Mon Dec 5 01:58:16 2016, gautam, Update, IMC, IMC ringdowns IMCringdown.pdfIMCringdown_2.pdf
          Reply  Mon Dec 5 15:55:25 2016, gautam, Update, IMC, IMC ringdowns 7x
             Reply  Mon Dec 5 19:29:52 2016, gautam, Update, IMC, IMC ringdowns REFLanomaly.pdf
                Reply  Tue Dec 6 00:43:41 2016, gautam, Update, IMC, more IMC ringdowns 
                   Reply  Wed Dec 7 11:52:48 2016, ericq, Update, IMC, Partial IMC ringdowns IMCpartial.pngIMCpartial.zip
                      Reply  Thu Dec 8 19:01:21 2016, rana, Update, IMC, Partial IMC ringdowns 
                   Reply  Sun Dec 2 17:26:58 2018, gautam, Update, IMC, IMC ringdown fitting tauTheoretical.pdfringdownFit.pdfringdownScatter.pdfcavPole.pdf
                      Reply  Fri Dec 7 12:51:06 2018, gautam, Update, IMC, IMC ringdown fitting time_reflAndTrans.pdfcorner_transOnly.pdfcorner_reflAndTrans.pdf
Message ID: 12667     Entry time: Tue Dec 6 00:43:41 2016     In reply to: 12666     Reply to this: 12672   14328
Author: gautam 
Type: Update 
Category: IMC 
Subject: more IMC ringdowns 

In an effor to see if I could narrow down the cause of the 100kHz ringing seen in the reflected PD signal, I tried a few things.

  1. Changed the PD - there was a PDA 255 sitting on the PSL table by the RefCav. Since it wasn't being used, I swapped the PD I was using with this. Unfortunately, this did not solve the problem.
  2. Used a different channel on the oscilloscope - ringing persisted
  3. Changed BNC cable running from PD to oscilloscope - ringing persisted
  4. Checked the spectrum of the PD under dark and steady illumination conditions for any features at 100kHz, saw nothing (as expected) 

I was working under the hypothesis that the ringing was due to some impedance mismatch between the PD output and the oscilloscope, and 4 above supports this. However, most documents I can find online, for example this one, recommend connecting the PD output via 50ohm BNC to a scope with input impedance 50ohms to avoid ringing, which is what I have done. But perhaps I am missing something.

Moreover, the ringdown in reflection actually supplies two of the five variables needed to apply the MIT method of loss estimation. I suppose we could fit the parameter "m4" from the ringdown in transmission, and then use this fitted value on the ringdown in reflection to see where the reflected power settles (i.e. the parameter "m3" as per the MIT paper). I will try analyzing the data on this basis.

I also measured the power levels at each of the PDs, these should allow us to calibrate the PD voltage outputs to power in Watts. All readings were taken with the Ophir power meter, with the filter removed, and the IMC locked.

PD Power level
REFL 0.47 mW (measured before 1.0 ND filter)
Trans 203 uW
Incident 1.06 mW

 

ELOG V3.1.3-